OpenStudy (anonymous):

Find the inverse Laplace transform of \[F(s)=\frac{3se^{-5s}}{s^2-8s+20}\]

OpenStudy (mendicant_bias):

Yo, you still hurr bb?

OpenStudy (mendicant_bias):

(Don't ask me why I said that, I'm just.....yeah.) Alright, this looks legitimately not fun. So, I can suggest some things immediately if you maybe haven't already thought of them. See the e^{as} term? What do you think will be relevant in the transform because of that term?

OpenStudy (mendicant_bias):

(I actually just took a test on totally relevant material, and I made a mistake specifically on a problem like this. I figured out afterwards.) >:C

OpenStudy (anonymous):

\[F(s)=\left(e^{-5s}\frac{3s}{s^2-8s+20}\right)\] Then whats the next step coz am stuck right here and confused

OpenStudy (anonymous):

Am not understanding this chapter at all.

OpenStudy (mendicant_bias):

Yeah, man, I had a hell of a time with it, too. That's a step in the right direction. I always get them confused in effect, but separating that e^{as} term will allow you to invoke either the Dirac Delta function, or the Unit Step/Heaviside function.

OpenStudy (mendicant_bias):

Alright, so that was a good first step. Now, you should consider how to treat the rest of the problem. Consider your remaining fraction as a standalone problem; do you have an idea of how you might solve it?

OpenStudy (anonymous):

\[\frac{ 3s }{ (s-4)^2+4 }\] I completed the square. then am not sure if it is \[e^{at} sin bt\] or \[e^{at} cos bt\]

OpenStudy (mendicant_bias):

There's no real mnemonic or memory tool I choose to remember it by, but the way I remember it as is that the one with the constant in the numerator is *always* the laplace of the sine function. The one with s in the numerator is *always* the laplace of the cosine function.

OpenStudy (mendicant_bias):

(Sorry for all but abandoning you on this problem for a while; I had some stuff going on and I got distracted.) Let me know when you come back if you still need help on this.

OpenStudy (anonymous):

@Mendicant_Bias I still need your help

OpenStudy (mendicant_bias):

Alright, the next time we're both on at the same time, do you understand which trig function it would be?

OpenStudy (anonymous):

i havent gotten anywhere on this problem. any hints?